# Byzantine Consensus in the Random Asynchronous Model

### v1

@@ 4. $n=2f+1$: Deterministic termination, safety whp
The main intuition behind this protocol is that, if a correct process accepts a value $v$, then all correct processes will accept $v$ by the end of the execution, and thus all correct processes will have the same set of accepted processes. Therefore, all correct processes can use a deterministic rule to decide the same value.
@@
10:            Receive a $(V_{j},p,r)$ $\triangleright$ receive single message per process, phase and round
11:            if $V_{j}$ contains a value $v$ with valid signatures from $phase$ distinct processes and $V_{i}[\textsc{Origin}(v)]=\varnothing$
@@
Proof.
We start by fixing two correct processes $p$ and $q$ and upper bounding the probability that $q$ does not receive any message from $p$ after $R$ iterations. The probability of not delivering $p$’s round-$1$ message to $q$ is at most $(1-\mathcal{C}(n,f))$ at each time step. There must be at least $R(n-f)^{2}$ time steps for $q$ to complete $R$ iterations, so the probability of $q$ not observing $p$’s message is at most
$(1-\mathcal{C}(n,f))^{R(n-f)^{2}}\approx e^{-R\mathcal{C}(n,f)(n-f)^{2}}.$
To finish the proof, we upper-bound the probability of any correct process not observing the input value of some other correct process. We first compute the expected number of (ordered) pairs of processes $(p,q)$ such that $q$ does not observe the input value of $p$ at the end of the $R$ iterations. There are $n(n-1)$ possible pairs, so this expected value is at most $E=n(n-1)e^{-R\mathcal{C}(n,f)(n-f)^{2}}$. Now, by Markov’s inequality, we have that
Pr(at least one unreachable $\Pr(\text{at least one unreachable pair})\leq E=n(n-1)e^{-R\mathcal{C}(n,f)(n-f)^{2}}.$
For fixed $n$ and $f$, this probability approaches 00 exponentially in the number of iterations $R$. ∎
@@
Proof.
By Lemma 4.1, every correct process receives at least one message from every other correct process in the first phase, whp. Since messages from correct processes always contain their input values with a valid signature, every correct processes $i$ will accept the input value of another correct process $j$ when $i$ receives the first message from $j$. ∎
@@
With $n=2f+1$, Algorithm 3 satisfies strong validity whp.
Proof.
Assume that all correct processes propose the same value $v$. Then, by Lemma 4.2, all correct processes will accept at least $n-f=f+1$ $v$ values whp. Since $f+1$ is a majority out of a maximum of $n=2f+1$ accepted values, correct processes decide $v$ whp. ∎
@@
If a value $v$ is accepted by a correct process $i$ at phase $p\leq f$, then all correct processes accept $v$ by the end of phase $p+1$, whp.
Proof.
If $i$ accepts $v$ at phase $p$, then $v$ must have signatures from at least $p$ processes. Process $i$ is not one of the $p$ processes, otherwise $i$ would have accepted $v$ at an earlier phase. Since $i$ accepts $v$, $i$ adds its signature to $v$ and will send $v$, as part of $V_{i}$, to all processes in phase $p+1$. By Lemma 4.1, all correct processes will thus receive $v$ by the end of phase $p+1$ whp, and will accept $v$ (if they haven’t already), since $v$ has the required number of signatures. ∎
@@
If a value $v$ is accepted by a correct process $i$ at phase $f+1$, then all correct processes accept $v$ by the end of phase $f+1$, whp.
Proof.
@@
So $j$ must have accepted $v$ at an earlier phase $p\leq f$ and therefore, by Lemma 4.4, all correct processes will accept $v$ by the end of phase $f+1$ whp. ∎
@@
With $n=2f+1$, Algorithm 3 satisfies agreement whp.
Proof.
By Lemma 4.4 and Lemma 4.5, with high probability, correct processes have the same set of accepted values by the end of phase $f+1$, and thus decide the same value. ∎
@@
Proof.
Follows immediately from the algorithm: correct processes only execute for $R(f+1)$ rounds. In each round, a correct process waits to receive $n-f$ messages from that round, which is guaranteed to occur since there are at least $n-f$ correct processes and the network is reliable (no-loss property). ∎
5. $n=f+2$: Deterministic Termination and Weak Validity, Agreement Whp
Interestingly, if we pose $n=f+2$, we can solve weak Byzantine consensus with deterministic validity and termination, and agreement whp, using the same protocol in Algorithm 3.
@@ 5. $n=f+2$: Deterministic Termination and Weak Validity, Agreement Whp
Proof.
If all processes are correct and propose the same value $v$, then all $(V_{i},phase,round)$ messages will have $v$ as their value, so no process can decide any other value. ∎
@@
With $n=f+2$, Algorithm 3 satisfies agreement whp.
Proof.
Lemmas 4.1, 4.2, 4.4, and 4.5 still hold: their proofs are also valid if $n=f+2$. Thus the proof of this theorem is the same as the proof of Theorem 4.6: By Lemma 4.4 and Lemma 4.5, with high probability, correct processes have the same set of accepted values by the end of phase $f+1$, and thus decide the same value. ∎
@@
Proof.
Correct processes only execute for $R(f+1)$ rounds. In each phase and round, a correct process waits to receive $n-f$ valid messages from that phase and round. This wait is guaranteed to terminate since there are at least $n-f$ correct processes, correct processes can always produce a valid message (a message $(V_{i},p,r)$ is valid if $p$ and $r$ are equal to the current phase and round, respectively), and the network is reliable (no-loss property). ∎
6. Negative Results
In this section we provide negative results that closely match our positive results from previous sections. Intuitively, we show that in the random asynchronous model it is not possible to obtain Byzantine consensus protocols with more powerful guarantees than the protocols we propose in this paper. This shows that the random asynchronous model, while avoiding some restrictions and impossibilities of the standard asynchronous model, is not overly permissive.
We prove the following three results. The first two results hold without any additional assumptions, while for the third result we require a couple of mild assumptions.
@@ 6. Negative Results
No protocol can solve Byzantine consensus in the random asynchronous model with deterministic strong validity, agreement, and termination.
@@
Before stating the third result, we describe two additional assumptions. Given a protocol $\mathcal{A}$, we say that an execution $E$ is clean if all processes are correct in $E$; a schedule $S$ is clean if at least one clean execution admits $S$ as its schedule. Our first assumption is that the total probability of clean schedules is not negligible. This assumption is reasonable, as without it, a protocol could, for instance, break agreement or validity in all clean executions and still claim whp agreement and validity; such a protocol would be useless in practice, where clean executions are common. A protocol satisfying this assumption is called clean.
Our second assumption is: if some correct process $p$ proposes $v\in\{0,1\}$ in an execution $E$ with schedule $S$, then there exists an execution $E^{\prime}$ with the same schedule $S$, in which $p$ proposes $1-v$. Intuitively, this assumption implies that processes can have the same communication pattern (e.g. send the same number of messages) whether they propose 00 or $1$. A protocol satisfying this assumption is called regular. Now we are ready to state our third negative result.
@@
With $n=f+2,f\geq 2$, no clean, regular protocol for Byzantine consensus in the random asynchronous model can ensure strong validity and agreement with high probability, while ensuring deterministic termination.
Proof sketch for Theorem 6.1.
@@
This result is equivalent to the FLP impossibility (Fischer et al., 1983) in the random asynchronous model, and the FLP proof holds in our model as well. Essentially, if at least one process can fail by crashing, there exists an infinite bivalent execution (the same execution as constructed in the FLP proof), which prevents processes from deciding without breaking agreement. ∎
Proof of Theorem 6.2.
We prove the result by contradiction. Assume that such a protocol $\mathcal{A}$ exists. We show through a “split-brain” argument that $\mathcal{A}$ admits an execution $E$ such that $E$ breaks agreement and $E$ has non-zero probability.
Take $E$ to be an execution in which at least one correct process proposes 00 and at least one correct process proposes $1$. Let $S_{0}$ be the set of correct processes that propose 00, $|S_{0}|\geq 1$, and $S_{1}$ be the set of correct processes that propose $1$, $|S_{1}|\geq 1$. Let $B$ be the set of Byzantine processes, $|B|=f$.
In $E$, the Byzantine processes behave toward the correct processes in $S_{0}$ as correct processes whose inputs are all 00. And they behave toward the $S_{1}$ as correct processes whose inputs are all $1$. Furthermore, any message between a process $S_{0}$ and a process in $S_{1}$, or vice-versa, is delayed by the scheduler until after all correct processes have decided.  [^2] Correct processes in $S_{0}$ cannot wait for messages from $S_{1}$ before proceeding with their protocol logic, as in asynchrony processes may only wait for messages from $n-f=f+1$ processes. In $E$, the $f+1$ processes that a process from $S_{0}$ hears from first always happen to be from $S_{0}$ or $B$. Symmetrically, processes from $S_{1}$ always hear from $S_{1}$ or $B$ first and cannot wait for messages from $S_{0}$. To processes in $S_{0}$, $E$ is indistinguishable from an execution in which all correct processes have input 00; thus they must decide 00 in a finite number of steps to preserve termination and strong validity. Symmetrically, to processes in $S_{0}$, $E$ is indistinguishable from an execution in which all correct processes have input $1$; thus they must decide $1$ in a finite number of steps. Therefore, $E$ is a finite execution in which agreement is violated.
Since $E$ is a finite execution, it has a finite number of scheduling steps, each of which has non-zero probability; thus, $E$ has non-zero probability. We have shown that $\mathcal{A}$ admits an execution that breaks agreement and has non-zero probability; a contradiction. ∎
Lemma 6.4.
With $n=f+2$, $f\geq 2$, let $\mathcal{A}$ be a clean, regular Byzantine consensus protocol which satisfies strong valdity and agreement whp, as well as deterministic termination. Then any clean schedule $S$ of $\mathcal{A}$ is bad for strong validity or agreement.
Proof.
@@
Consider any clean schedule $S$ of $\mathcal{A}$. Since $\mathcal{A}$ is regular, there exists a clean execution $E$ of $\mathcal{A}$ with schedule $S$, in which at least 2 processes propose 00 and at least 2 processes propose $1$. Assume wlog that $p_{1}$ and $p_{2}$ propose 00, while $p_{n-1}$ and $p_{n}$ propose $1$. We now describe three executions $E_{1}$, $E_{2}$, and $E_{3}$, with the same schedule $S$ as $E$, and show that $\mathcal{A}$ must break either strong validity or agreement in one of the three executions. This is sufficient to show that $S$ is bad for strong validity or agreement.
@@
Let $E_{1}$ be an execution with schedule $S$, in which all processes behave identically to $E$; in $E_{1}$, $p_{1}$ and $p_{2}$ are correct, while $p_{3},\ldots,p_{n}$ are Byzantine but behave correctly. To satisfy strong validity, $p_{1}$ must decide 00 in $E_{1}$.
@@
Let $E_{2}$ be an execution with schedule $S$, in which again all processes behave identically to $E$; this time, $p_{1}$ and $p_{n}$ are correct, while $p_{2},\ldots,p_{n-1}$ are Byzantine. Since $E_{1}$ and $E_{2}$ are indistinguishable to $p_{1}$, and $p_{1}$ has deterministic logic, $p_{1}$ must decide the same value in both executions, namely 00. Thus, in order to satisfy agreement, $p_{n}$ must also decide 00 in $E_{2}$.
@@
Let $E_{3}$ be an execution with schedule $S$, in which again all processes behave as in $E$; this time, $p_{n-1}$ and $p_{n}$ are correct, while $p_{1},\ldots,p_{n-2}$ are Byzantine. Since $E_{2}$ and $E_{3}$ are indistinguishable to $p_{n}$, and $p_{n}$ has deterministic logic, $p_{n}$ must decide the same value in both executions, namely 00. But this breaks strong validity, as both correct processes ($p_{n-1}$ and $p_{n}$) have proposed $1$ in $E_{3}$. ∎
Proof of Theorem 6.3.
Assume by contradiction that such a protocol $\mathcal{A}$ exists. By Lemma 6.4, all clean schedules of $\mathcal{A}$ are bad for strong validity or agreement, and since $\mathcal{A}$ is clean, the total probability of its clean schedules is not negligible. It follows that the total probability of schedules which are bad for strong validity or agreement is not negligible, so $\mathcal{A}$ cannot satisfy strong validity and agreement both with high probability; a contradiction. ∎
7. Related Work
Random scheduling

### v2

@@ 4 $n=2f+1$: Deterministic Termination, Safety with High Probability
The main intuition behind this protocol is that, if a correct process accepts a value $v$, then all correct processes will accept $v$ by the end of the execution, and thus all correct processes will have the same set of accepted values. Therefore, all correct processes can use a deterministic rule to decide the same value.
@@
10:            Receive a $(V_{j},p,r)$ $\triangleright$ receive single msg per process, phase and round
11:            if $\exists v\in V_{j}:$ $v$ is signed by $phase$ distinct processes and $V_{i}[\textsc{Origin}(v)]=\varnothing$
@@
Proof 4.2.
We start by fixing two correct processes $p$ and $q$ and upper bounding the probability that $q$ does not receive any message from $p$ after $R$ iterations. The probability of not delivering $p$’s round-$1$ message to $q$ is at most $(1-\mathcal{C}(n,f))$ at each time step. There must be at least $R(n-f)$ time steps for $q$ to complete $R$ iterations, since $q$ waits for at least $n-f$ messages at each iteration, and each message received consumes one time step. Thus, using assumptions (1) and (2), the probability of $q$ not observing $p$’s message is at most
$(1-\mathcal{C}(n,f))^{R(n-f)}\approx e^{-R\mathcal{C}(n,f)(n-f)}.$
To finish the proof, we upper-bound the probability of any correct process not observing the input value of some other correct process. We first compute the expected number of (ordered) pairs of processes $(p,q)$ such that $q$ does not observe the input value of $p$ at the end of the $R$ iterations. There are $n(n-1)$ possible pairs, so this expected value is at most $E=n(n-1)e^{-R\mathcal{C}(n,f)(n-f)}$. Now, by Markov’s inequality, we have that
Pr(at least one unreachable $\Pr(\text{at least one unreachable pair})\leq E=n(n-1)e^{-R\mathcal{C}(n,f)(n-f)}.$
For fixed $n$ and $f$, this probability approaches 00 exponentially in $R$.
@@
Proof 4.4.
By Lemma 4.1, every correct process receives at least one message from every other correct process in the first phase, w.h.p. Since messages from correct processes always contain their input values with a valid signature, every correct processes $i$ will accept the input value of another correct process $j$ when $i$ receives the first message from $j$.
@@
With $n=2f+1$, Algorithm 3 satisfies strong validity w.h.p.
Proof 4.6.
Assume that all correct processes propose the same value $v$. Then, by Lemma 4.3, all correct processes will accept at least $n-f=f+1$ copies of value $v$ w.h.p. Since $f+1$ is a majority out of a maximum of $n=2f+1$ accepted values, correct processes decide $v$ w.h.p.
@@
If a value $v$ is accepted by a correct process $i$ at phase $p\leq f$, then all correct processes accept $v$ by the end of phase $p+1$, w.h.p.
Proof 4.8.
If $i$ accepts $v$ at phase $p$, then $v$ must have signatures from at least $p$ processes. Process $i$ is not one of the $p$ processes, otherwise $i$ would have accepted $v$ at an earlier phase. Since $i$ accepts $v$, $i$ adds its signature to $v$ and will send $v$, as part of $V_{i}$, to all processes in phase $p+1$. By Lemma 4.1, all correct processes will thus receive $v$ by the end of phase $p+1$ w.h.p., and will accept $v$ (if they haven’t already), since $v$ has the required number of signatures.
@@
If a value $v$ is accepted by a correct process $i$ at phase $f+1$, then all correct processes accept $v$ by the end of phase $f+1$, w.h.p.
Proof 4.10.
@@
So $j$ must have accepted $v$ at an earlier phase $p\leq f$ and therefore, by Lemma 4.7, all correct processes will accept $v$ by the end of phase $f+1$ w.h.p.
@@
With $n=2f+1$, Algorithm 3 satisfies agreement w.h.p.
Proof 4.12.
By Lemma 4.7 and Lemma 4.9, with high probability, correct processes have the same set of accepted values by the end of phase $f+1$, and thus decide the same value.
@@
Proof 4.14.
Follows immediately from the algorithm: correct processes only execute for $R(f+1)$ rounds. In each round, a correct process waits to receive $n-f$ messages from that round, which is guaranteed to occur in a finite number of steps, since there are at least $n-f$ correct processes and the network is reliable (no-loss property).
5 $n=f+2$: Deterministic Termination and Validity, Agreement w.h.p.
Interestingly, for $n=f+2$, we can solve weak Byzantine consensus with deterministic validity and termination, and agreement w.h.p., using the same protocol in Algorithm 3.
@@
Proof 5.2.
If all processes are correct and propose the same value $v$, then all $(V_{i},phase,round)$ messages will have $v$ as their value, so no process can decide any other value.
@@
With $n=f+2$, Algorithm 3 satisfies agreement w.h.p.
Proof 5.4.
Lemmas 4.1, 4.3, 4.7, and 4.9 still hold: their proofs are also valid if $n=f+2$. Thus the proof of this theorem is the same as the proof of Theorem 4.11: By Lemma 4.7 and Lemma 4.9, with high probability, correct processes have the same set of accepted values by the end of phase $f+1$, and thus decide the same value.
@@
Proof 5.6.
Correct processes only execute for $R(f+1)$ rounds. In each phase and round, a correct process waits to receive $n-f$ valid messages from that phase and round. This wait is guaranteed to terminate since there are at least $n-f$ correct processes, correct processes can always produce a valid message (a message $(V_{i},p,r)$ is valid if $p$ and $r$ are equal to the current phase and round, respectively), and the network is reliable (no-loss property).
6 Negative Results
In this section we provide negative results that closely match our positive results from previous sections. Intuitively, we show that in the random asynchronous model it is not possible to obtain Byzantine consensus protocols with more powerful guarantees than the protocols we propose in this paper. This shows that the random asynchronous model, while avoiding some restrictions and impossibilities of the standard asynchronous model, is not overly permissive. Concretely, we prove the following three results.
@@ 6 Negative Results
No protocol can solve consensus in the random asynchronous model with deterministic strong validity, agreement, and termination, if at least one process may fail by crashing.
@@
With $n=f+2,f\geq 2$, no protocol for Byzantine consensus in the random asynchronous model can ensure strong validity and agreement with high probability, while ensuring deterministic termination.
Proof 6.4 (Proof sketch for Theorem 6.1).
@@
This result is equivalent to the FLP impossibility [13] in the random asynchronous model, and the FLP proof holds in our model as well. Essentially, if at least one process can fail by crashing, there exists an infinite bivalent execution (the same execution as constructed in the FLP proof), which prevents processes from deciding without breaking agreement.
Instead of proving Theorem 6.2 directly, we can prove the following stronger result, which also implies Theorem 6.2.
Theorem 6.5.
No protocol for consensus in the random asynchronous model can ensure strong validity and agreement almost surely, as well as deterministic termination, if at least one process may fail by crashing.
Proof 6.6.
Assume such a protocol $\mathcal{A}$ exists. Since $\mathcal{A}$ ensures deterministic termination, it must terminate in finite time in every execution. Furthermore, $\mathcal{A}$ ensures strong validity and agreement almost surely, so the probability of schedules in which $\mathcal{A}$ terminates in a bivalent state (in the sense of FLP [13]) must be 00. Yet, following the FLP proof, there exists an infinite bivalent execution $E$. It follows that $\mathcal{A}$ must decide in a finite prefix $\pi$ of $E$, when the state is still bivalent. Since $\pi$ has finite length, its schedule has non-zero probability. We have shown that $\mathcal{A}$ does not ensure agreement and validity almost surely, a contradiction.
Proof 6.7 (Proof of Theorem 6.3).
Assume towards a contradiction that such a protocol $\mathcal{A}$ exists. Call an execution $E$ of $\mathcal{A}$ heterogeneous if in $E$ at least 2 processes propose 00 and at least 2 processes propose $1$. Assume wlog that $p_{1}$ and $p_{2}$ propose 00, while $p_{n-1}$ and $p_{n}$ propose $1$.
Let $E$ be a heterogeneous execution in which all processes are correct. Let $S$ be the schedule of $E$. Let $\pi$ be shortest prefix of $S$ in which all processes have decided; in other words, we truncate any steps in $S$ after the processes have decided. The prefix $\pi$ must be finite since $\mathcal{A}$ ensures deterministic termination.
@@
We now describe three executions $E_{1}$, $E_{2}$, and $E_{3}$, with the same schedule $\pi$, and show that $\mathcal{A}$ must break either strong validity or agreement in one of the executions. This is sufficient to show that $\pi$ is bad (in the sense of Section 2) for strong validity or agreement.
@@
Let $E_{1}$ be an execution with schedule $\pi$, in which all processes behave identically to $E$; in $E_{1}$, $p_{1}$ and $p_{2}$ are correct, while $p_{3},\ldots,p_{n}$ are Byzantine but behave correctly. By strong validity, $p_{1}$ must decide 00 in $E_{1}$.
@@
Let $E_{2}$ be an execution with schedule $\pi$, in which again all processes behave identically to $E$; this time, $p_{1}$ and $p_{n}$ are correct, while $p_{2},\ldots,p_{n-1}$ are Byzantine. Since $E_{1}$ and $E_{2}$ are indistinguishable to $p_{1}$, and $p_{1}$ has deterministic logic, $p_{1}$ must decide the same value in both executions, namely 00. Thus, in order to satisfy agreement, $p_{n}$ must also decide 00 in $E_{2}$.
@@
Let $E_{3}$ be an execution with schedule $\pi$, in which again all processes behave as in $E$; this time, $p_{n-1}$ and $p_{n}$ are correct, while $p_{1},\ldots,p_{n-2}$ are Byzantine. Since $E_{2}$ and $E_{3}$ are indistinguishable to $p_{n}$, and $p_{n}$ has deterministic logic, $p_{n}$ must decide the same value in both executions, namely 00. But this breaks strong validity, as both correct processes ($p_{n-1}$ and $p_{n}$) have proposed $1$ in $E_{3}$.
We have shown that $\pi$ is bad for agreement or strong validity. Since the length of $\pi$ is fixed (and finite) for fixed $n$ and $f$, the probability of $\pi$ is non-negligible according to our definition in Section 2. Therefore, $\mathcal{A}$ does not ensure strong validity and agreement with high probability, a contradiction.
7 The Random Asynchronous Model vs Standard Models
In this section we compare the random asynchronous model with the standard network models—asynchrony, synchrony, and partial synchrony—in terms of task solvability. We show that it is (i) strictly stronger than full asynchrony (i.e., the set of tasks solvable in random asynchrony is a strict superset of the set of tasks solvable in asynchrony), (ii) strictly weaker than synchrony, and (iii) incomparable with partial synchrony, with respect to task solvability. Table 2 summarizes the relationships.
Table 2: Relative power of the random asynchronous model.
Asynchrony Partial synchrony Synchrony
Random asynchrony strictly stronger incomparable strictly weaker
Asynchrony. Every task solvable in asynchrony is also solvable in random asynchrony because every finite prefix of an asynchronous schedule occurs with non-zero probability under our scheduler. [^2] Formally, let $S$ be any (possibly infinite) asynchronous schedule and $\pi$ a finite prefix of $S$. Since each delivery step is chosen independently with lower-bound probability $C(n,f)\!>\!0$, the probability of $\pi$ is at least $C(n,f)^{|\pi|}$. Conversely, the task “Byzantine binary consensus with $n=2f+1$, deterministic termination, and safety holding with high probability” is solvable in the random asynchronous model (Section 4) yet impossible in pure asynchrony by the standard network partition argument [7, 8]. Hence the random asynchronous model is strictly stronger.
Synchrony. The random asynchronous model is strictly more restrictive than synchrony: any task that is solvable in the random asynchronous model is also solvable in synchrony, and there are tasks which are solvable in synchrony but not in the random asynchronous model. As an example of the latter, consider deterministic crash-fault tolerant consensus: this task is impossible in the random asynchronous model (Theorem 6.1) but it is solvable in synchrony.
We now prove the former direction: any task solvable in the random asynchronous model is also solvable in synchrony. Take any task $\tau$ and let $\mathcal{A}_{RA}$ be a protocol that solves $\tau$ in the random asynchronous model. We can simulate $\mathcal{A}_{RA}$ in synchrony using an adapter $\Sigma_{p}$ for each process $p$ between $\mathcal{A}_{RA}$ and the network, in the following way:
  1. 1.
    At each synchronous time step $t=0,1,\ldots$, each correct process $p$ executes all outstanding local steps of $\mathcal{A}_{RA}$, including those triggered by messages delivered since the previous time step. Process $p$ sends the same messages it would in $\mathcal{A}_{RA}$, by handing them to $\Sigma_{p}$.
  2. 2.
    The synchronous network delivers all messages send at time $t$ by time $t+1$. $\Sigma_{p}$ buffers any messages received by $p$ between $t$ and $t+1$.
  3. 3.
    At time $t+1$, $\Sigma_{p}$ draws a fresh random permutation of the links on which $p$ has received messages, then delivers to $\mathcal{A}_{RA}$ the earliest pending message on each link in that order.
The resulting schedule visible to $\mathcal{A}_{RA}$ is valid in the random asynchronous model, as any causal relation is preserved. Since the $\Sigma_{p}$ adapters collectively deliver messages from at most $n^{2}$ links at each synchronous time step, each link has a probability of at least $1/n^{2}:=\mathcal{C}(n,f)$ of being selected at each time step internal to the random asynchronous protocol $\mathcal{A}_{RA}$. Thus, the $\Sigma_{p}$ adapters preserve the probabilistic delivery properties of the random asynchronous model. We have described a construction that simulates any random asynchronous algorithm in synchrony, thus showing that any task that is solvable in the random asynchronous model is solvable in synchrony.
Partial synchrony. The random asynchronous model is incomparable with respect to partial synchrony in terms of solvability: there are tasks that are solvable in the random asynchronous model, but not in partial synchrony, and vice-versa.
As an example of a task that is solvable in the random asynchronous model but not in partial synchrony, recall the same $n=2f+1$ Byzantine consensus task above. Section 4 shows that this task is solvable in the random asynchronous model. However, this task is not solvable in partial synchrony [12].
To show that some tasks are solvable in partial synchrony but not in the random asynchronous model, consider the task of (deterministic) crash-fault tolerant consensus. This task is solvable in partial synchrony if $n\geq 2f+1$ [9], but is not solvable in the random asynchronous model if at least one process can crash (Theorem 6.1).
Discussion. Although the random–asynchronous model is strictly weaker than full synchrony in terms of task solvability, it can approximate synchronous behavior with high probability. After $R$ rounds of all-to-all communication, every pair of correct processes has exchanged at least one message with probability $1-\exp(-\Omega(R))$ (Lemma 4.1). This mirrors the deterministic guarantee offered by a single synchronous round—modulo the known time bound $\Delta$.
Leveraging this property, we can implement, in the random asynchronous model, a crash-fault detector $\sim\!\mathcal{P}$ that is perfect with high probability:
Strong completeness
Every process that crashes is eventually permanently suspected by every correct process.
Strong accuracy w.h.p.
No correct process is suspected by any correct process w.h.p.
The implementation is a classic heartbeat scheme. Each process broadcasts a heartbeat message once per round; process $p$ suspects $q$ iff it has not received a heartbeat from $q$ in the last $R$ rounds. As we increase $R$, the probability of false positive suspicion decreases exponentially, through an argument similar to Lemma 4.1.
Such a failure detector is similar to the perfect $\mathcal{P}$ and eventually perfect $\diamond\mathcal{P}$ failure detectors, which can be implemented in synchrony and partial synchrony respectively [9]. Whether protocols that rely on $\mathcal{P}$ or $\Diamond\mathcal{P}$ can be systematically translated to use $\sim\!\mathcal{P}$—preserving their guarantees w.h.p.—remains an open question.
Intuitively, such a systematic translation seems unlikely to exist, as downgrading deterministic communication guarantees to hold w.h.p. could break arbitrary internal algorithm invariants. Consider a synchronous algorithm that runs for $T$ rounds; such an algorithm is able to rely on the fact that it will receive messages from the same set of at least $n-f$ correct processes in each of the $T$ rounds—e.g., at the end of round $T$ each process $p$ may output the $(n-f)^{\text{th}}$ highest process id $p$ heard from in every one of the $T$ rounds. However, if we directly translate this algorithm to random asynchrony, by replacing each synchronous round with $R$ asynchronous rounds (for at total of $RT$ rounds), the guarantee of receiving messages from a common subset of $n-f$ processes holds merely with high probability, not deterministically. In the rare executions where $p$ hears from a common subset of fewer than $n-f$ processes, the translated algorithm would try to access data that does not exist, violating its invariants. Any translation from synchrony to random asynchrony would therefore need manual, algorithm-specific fallback logic to cope with those low-probability outliers, making a generic translation unlikely.
8 Related Work